OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6> can you please explain step by step?

OpenStudy (anonymous):

you are to find the Dot Product of \(v\) and \(w\)..

OpenStudy (anonymous):

Just multiply the corresponding elements of the vector \(v\) and \(w\), and sum them..

OpenStudy (anonymous):

\((3 \times -4) + (-8 \times -2) + (-3 \times -6) = ??\)

OpenStudy (anonymous):

ohhh okay! haha thanks :)

OpenStudy (anonymous):

If you come across any doubt, then let me know.. :)

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

do you know how to do this kind: Let u = <-5, -9>, v = <6, 8>. Find -8u - 2v. ?

OpenStudy (anonymous):

Firstly multiply -8 to all elements of \(u\)..

OpenStudy (anonymous):

Then multiply \(-2\) with all elements of \(v\).. Then simply add them..

OpenStudy (anonymous):

40 + 72 -12 + (-16) like that?

OpenStudy (anonymous):

you can only add or subtract corresponding elements of the two vector.. First of one with first of second.. :)

OpenStudy (anonymous):

See like this, I show you...

OpenStudy (anonymous):

oh so 40 + (-12) 72 + (-16)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

But after adding, write result as <a,b>.. :)

OpenStudy (anonymous):

ohhh okay!

OpenStudy (anonymous):

Like : \(40 - 12 = 28\) and \(72-16 = 56\) Then it will be : \(<28, 56>\)

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (anonymous):

In vectors, always add corresponding elements of two or more vectors only..

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

Yeah, keep it up, the more you practise, the more you will be gaining.. :)

OpenStudy (anonymous):

thanks soooo much! that was really helpful, now I can do all of my problems :)

OpenStudy (anonymous):

Good... :)